# PrivaDE: Privacy-preserving Data Evaluation for Blockchain-based Data Marketplaces

### v2

@@ Abstract.
conference: ACM ASIA Conference on Computer and Communications Security; June 01–06, 2026; Bangalore, India
@@ 1. Introduction
Figure 1. Overall design of PrivaDE. The core protocol consists of two parts: a secure inference of the data points and a scoring function that takes the inference results and labeled data points as input. Several optimization components, detailed in Section 5, are applied before secure inference to reduce the overhead of the first component.
@@ 4.3. The PrivaDE Protocol
  • $(\mathsf{SetupCom},\mathsf{Commit},\mathsf{Open})$ is a commitment scheme satisfying Definition 2.2.1.
@@
    $L_{\textsf{ML-Inf}}=\left\{\,\begin{array}[]{l}(\mathsf{pp},{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\overrightarrow{x},\mathtt{com}_{\theta},\overrightarrow{\mathtt{com}_{y^{\prime}}}},\\ {\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\theta,\overrightarrow{\mathtt{com}_{x}},\overrightarrow{y^{\prime}}})\end{array}:\begin{array}[]{l}\exists\,({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\theta,r_{\theta},\overrightarrow{y^{\prime}},\overrightarrow{r_{y^{\prime}}}},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overrightarrow{x},\overrightarrow{r_{x}}})\ \text{s.t.}\\ M_{\theta}(x)=y^{\prime},\\ {\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathsf{Open}(\mathsf{pp},\mathtt{com}_{\theta},\theta,r_{\theta})=1},\\ {\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\forall y^{\prime}\,\mathsf{Open}(\mathsf{pp},\mathtt{com}_{y^{\prime}},y^{\prime},r_{y^{\prime}})=1},\\ {\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\forall x\,\mathsf{Open}(\mathsf{pp},\mathtt{com}_{x},x,r_{x})=1}\end{array}\right\}.$
@@ 5.2. Split Model
Figure 6. Conceptual diagram of split model in PrivaDE: Split A evaluated in a secure 2PC; split B evaluated by the data owner; split C evaluated by the model owner.
@@ 5.3. Representative Set Selection
which preserve distances up to $\varepsilon$-distortion by the Johnson–Lindenstrauss lemma (johnson1984extensions). Alice and Bob jointly sample $R$ via $md$ secure coin flips (blum1981coin), deriving a shared pseudorandom seed. Bob projects locally and runs representative-selection on $\{x^{\prime}\}$, reducing computation and communication while retaining fidelity.
@@ 6. Blockchain Integration for PrivaDE
Figure 7. Decentralized Data Marketplace Workflow with PrivaDE integration. Commitments and protocol results are posted on-chain for transparency, and data transaction is seamlessly integrated.
@@ 7. Evaluation
Our open-source implementation is publicly available. We instantiate SMPC using the SPDZ2k protocol within MP-SPDZ (mpspdz), employ EZKL (zkonduit_ezkl) for ZKP-based inference, and implement the challenge protocol in Circom (circom).
@@ 7.1.1. Scalability Test
Figure 8. Online runtime versus choices of $n$, $k$, $|I|$, and the verification ratio in $\mathsf{CnCZK}$. All parameters except $n$ exhibit approximately linear scaling.
@@ 7.2. Robustness of the scoring algorithm
Figure 9. Robustness of our multi-point scoring algorithm vs. other active learning methods. The results show our method is comparable to core-set and outperforms random and entropy sampling.
@@ A.2. The CnCZK Protocol
Notation.
Let $M_{\theta}$ be a fixed $L$-layer network with parameters $\theta$ (with $\theta_{\ell}$ the weights at layer $\ell$), inputs $\{x_{i}\}_{i=1}^{N}$, intermediate activations $a_{\ell,i}$ for $\ell\in\{0,\dots,L\}$ (with $a_{0,i}=x_{i}$), and outputs $y^{\prime}_{i}=a_{L,i}$. Denote the $\ell$-th layer by $m_{\ell}(\theta_{\ell};\cdot)$, so $a_{\ell,i}=m_{\ell}(\theta_{\ell};a_{\ell-1,i})$.
@@ Appendix D Security Proofs
    $\mathcal{S}_{\mathsf{Bob}}$ creates $k$ dummy commitments $(\mathtt{com}_{y^{\prime}_{i}},r_{y^{\prime}_{i}})\leftarrow\mathsf{Commit}(\mathsf{pp},0)$ for $i\in[k]$. It then invokes $\mathcal{S}_{\mathsf{ZK}}$ for $L_{\textsf{ML-Inf}}$ (hidden weights, public inputs) to generate $\pi\leftarrow\mathcal{S}_{\mathsf{ZK}}(\mathsf{pp},\overrightarrow{a_{B}},\mathtt{com}^{C}_{0},\{\mathtt{com}_{y^{\prime}_{i}}\}_{i})$. It sends $\pi,\mathtt{com}_{y^{\prime}_{1}},\ldots,\mathtt{com}_{y^{\prime}_{k}}$ to $\mathcal{A}$. If $\mathcal{A}$ aborts, it aborts.
@@
    $\mathcal{S}_{\mathsf{Bob}}$ invokes $\mathcal{A}$ to obtain $(D_{B}^{x},y_{1},\ldots,y_{k},r_{y_{1}},\ldots,r_{y_{k}},\mathtt{com}_{y^{\prime}_{1}},\ldots,\mathtt{com}_{y^{\prime}_{k}})$ for scoring.
@@ Appendix E Model Definitions
Here we show the exact model definitions of the distilled model in our experiment and the A-B-C splits. The split between B–C is evaluated with a data reconstruction attack, requiring the average SSIM score to be below 0.3.

### v3

@@ Abstract.
Privacy-preserving machine learning, Secure data valuation, Zero-knowledge proofs, Secure model inference, Active learning, Data marketplaces, Blockchain-based systems, Decentralized machine learning, Cryptographic protocols
journalyear: 2026copyright: ccconference: ACM Asia Conference on Computer and Communications Security; June 01–05, 2026; Bangalore, Indiabooktitle: ACM Asia Conference on Computer and Communications Security (ASIA CCS ’26), June 01–05, 2026, Bangalore, Indiadoi: 10.1145/3779208.3785393isbn: 979-8-4007-2356-8/2026/06ccs: Security and privacy Privacy-preserving protocolsccs: Computing methodologies Supervised learningccs: Security and privacy Security protocolsccs: Security and privacy Cryptographyccs: Information systems Data mining
@@ 1. Introduction
Figure 1. Overall design of PrivaDE. The core protocol consists of two parts: a secure inference of the data points and a scoring function that takes the inference results and labeled data points as input. Several optimization components, detailed in Section 5, are applied before secure inference to reduce the overhead of the first component.:
@@ 4.3. The PrivaDE Protocol
  • $(\mathsf{SetupCom},\mathsf{Commit},\mathsf{Open})$ is a commitment scheme satisfying Definition C.1.
@@
    $L_{\textsf{ML-Inf}}=\left\{\,\begin{array}[]{@{}l@{}}(\mathsf{pp},{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\overrightarrow{x},\mathtt{com}_{\theta},\overrightarrow{\mathtt{com}_{y^{\prime}}}},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\theta,\overrightarrow{\mathtt{com}_{x}},\overrightarrow{y^{\prime}}}):\\[1.99997pt] \exists\,({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\theta,r_{\theta},\overrightarrow{y^{\prime}},\overrightarrow{r_{y^{\prime}}}},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overrightarrow{x},\overrightarrow{r_{x}}})\ \text{s.t.}\\ M_{\theta}(x)=y^{\prime},\\ {\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathsf{Open}(\mathsf{pp},\mathtt{com}_{\theta},\theta,r_{\theta})=1},\\ {\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\forall y^{\prime}\,\mathsf{Open}(\mathsf{pp},\mathtt{com}_{y^{\prime}},y^{\prime},r_{y^{\prime}})=1},\\ {\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\forall x\,\mathsf{Open}(\mathsf{pp},\mathtt{com}_{x},x,r_{x})=1}\end{array}\right\}.$
@@ 5.2. Split Model
Figure 6. Conceptual diagram of split model in PrivaDE: Split A evaluated in a secure 2PC; split B evaluated by the data owner; split C evaluated by the model owner.:
@@ 5.3. Representative Set Selection
which preserve distances up to $\varepsilon$-distortion by the Johnson - Lindenstrauss lemma (johnson1984extensions). Alice and Bob jointly sample $R$ via $md$ secure coin flips (blum1981coin), deriving a shared pseudorandom seed. Bob projects locally and runs representative-selection on $\{x^{\prime}\}$, reducing computation and communication while retaining fidelity.
@@ 6. Blockchain Integration for PrivaDE
Figure 7. Decentralized Data Marketplace Workflow with PrivaDE integration. Commitments and protocol results are posted on-chain for transparency, and data transaction is seamlessly integrated.:
@@ 7. Evaluation
Our open-source implementation is publicly available (tpmmthomas_sdv). We instantiate SMPC using the SPDZ2k protocol within MP-SPDZ (mpspdz), employ EZKL (zkonduit_ezkl) for ZKP-based inference, and implement the challenge protocol in Circom (circom).
@@ 7.1.1. Scalability Test
Figure 8. Online runtime versus choices of $n$, $k$, $|I|$, and the verification ratio in $\mathsf{CnCZK}$. All parameters except $n$ exhibit approximately linear scaling.:
@@ 7.2. Robustness of the scoring algorithm
Figure 9. Robustness of our multi-point scoring algorithm vs. other active learning methods. The results show our method is comparable to core-set and outperforms random and entropy sampling.:
@@ A.2. The CnCZK Protocol
Notation. Let $M_{\theta}$ be a fixed $L$-layer network with parameters $\theta$ (with $\theta_{\ell}$ the weights at layer $\ell$), inputs $\{x_{i}\}_{i=1}^{N}$, intermediate activations $a_{\ell,i}$ for $\ell\in\{0,\dots,L\}$ (with $a_{0,i}=x_{i}$), and outputs $y^{\prime}_{i}=a_{L,i}$. Denote the $\ell$-th layer by $m_{\ell}(\theta_{\ell};\cdot)$, so $a_{\ell,i}=m_{\ell}(\theta_{\ell};a_{\ell-1,i})$.
@@ Appendix D Security Proofs
    $\mathcal{S}_{\mathsf{Bob}}$ creates $k$ dummy commitments $(\mathtt{com}_{y^{\prime}_{i}},r_{y^{\prime}_{i}})\leftarrow$
    $\mathsf{Commit}(\mathsf{pp},0)$ for $i\in[k]$. It then invokes $\mathcal{S}_{\mathsf{ZK}}$ for $L_{\textsf{ML-Inf}}$ (hidden weights, public inputs) to generate $\pi\leftarrow$
    $\mathcal{S}_{\mathsf{ZK}}(\mathsf{pp},\overrightarrow{a_{B}},\mathtt{com}^{C}_{0},\{\mathtt{com}_{y^{\prime}_{i}}\}_{i})$. It sends $\pi,\mathtt{com}_{y^{\prime}_{1}},\ldots,\mathtt{com}_{y^{\prime}_{k}}$ to $\mathcal{A}$. If $\mathcal{A}$ aborts, it aborts.
@@
    $\mathcal{S}_{\mathsf{Bob}}$ invokes $\mathcal{A}$ to obtain $(D_{B}^{x},y_{1},\ldots,y_{k},$ $r_{y_{1}},\ldots,r_{y_{k}},$
    $\mathtt{com}_{y^{\prime}_{1}},\ldots,\mathtt{com}_{y^{\prime}_{k}})$ for scoring.